# How Uniform Random Weights Induce Non-uniform Bias: Typical Interpolating Neural Networks Generalize with Narrow Teachers

### v1

@@ Appendix A Generalization Results
  • A restatement and discussion on the connection to a well-known result from [Shalev-Shwartz and Ben-David, 2014, Corollary 2.3] in Appendix A.1.
  • The proof of our primary generalization result (Theorem 3.2) in Appendix A.2.
  • The proof of Corollary 3.3 in Appendix A.3.
  • A further discussion on the relation of Theorem 3.2 to PAC-Bayes is brought in Appendix A.4
  • In Appendix A.5 we prove a refined version of Theorem 3.2 and discuss its implications.
  • The proof of an alternative generalization result, used for proving generalization in the continuous setting, in Appendix A.6.
A.1 Finite Hypothesis Class PAC Generalization Bound
Here we present a classic generalization bound for finite hypothesis class adapted from Shalev-Shwartz and Ben-David [2014], which we will use throughout our paper.
Theorem A.1.
[[[[adapted from [Shalev-Shwartz and Ben-David, 2014, Corollary 2.3]]normal-]]] Let $\mathcal{H}$ be a finite hypothesis class. Let $\delta\in\left(0,1\right)$ and $\epsilon>0$ and let $N$ be an integer that satisfies
$N\geq\frac{\log\left({\left\lvert{\mathcal{H}}\right\rvert}/\delta\right)}{\epsilon}\,.$
Then, for any realizable data distribution $\mathcal{D}$, with probability of at least $1-\delta$ over the choice of an i.i.d sample $S$ of size $N$, we have that for every interpolating101010[Shalev-Shwartz and Ben-David, 2014, Corollary 2.3] is more general, but we focus on interpolating hypotheses. hypothesis, $h_{\bm{\theta}}$, it holds that
$\mathcal{L}_{\mathcal{D}}\left(h_{\mathcal{S}}\right)\leq\epsilon\,.$
Applying Theorem A.1 to quantized neural networks. It is possible to directly obtain a generalization bound for any quantized model, such as those presented in Section 5. Note that the hypothesis class of quantized models with a finite number of parameters, for example, neural networks with finite width and depth, with $M$ parameters and $Q$ quantization levels is finite and of size $\left|\mathcal{H}\right|=Q^{M}$. Therefore, the sample complexity from Theorem A.1 becomes
$N\geq\frac{M\log\left(Q\right)+\log\left(\frac{1}{\delta}\right)}{\epsilon}$ (12)
Note that Theorem A.1 is used to prove Theorem 3.2.The sample complexity in (12) is dependent on the student parameters, and therefore much worse than the sample complexities we derive throughout our paper.
A.2 Proving the Generalization of Algorithm 1
In this section, we rely on the existence of the teacher to provide guarantee on the generalization of Algorithm 1. We will first restate Theorem 3.2:
Theorem A.2 (Theorem 3.2 restated).
Let $\epsilon\in\left(0,1\right)$ and $\delta\in\left(0,\frac{1}{5}\right)$, and assume that $\tilde{p}<\frac{1}{2}$. For any $N$ larger than
$\left(-\log\left({\tilde{p}}\right)+3\log\left(\frac{2}{\delta}\right)\right)\frac{1}{\epsilon}\,,$
@@ A.2 Proving the Generalization of Algorithm 1
$\mathbb{P}_{S\sim\mathcal{D}^{N},h\sim\mathcal{P}_{\mathcal{S}}}\left(\mathcal{L}_{\mathcal{D}}(h)\leq\epsilon\right)\geq 1-\delta\,.$
Proof Outline. The idea is to show that the probability to sample a TE model within a finite number of steps $\tau$ is large. Conditioning on that event, we treat those models as a realizable finite hypothesis class, obtaining the sample complexity using Theorem A.1, which is then bounded, and after some technical details the theorem is obtained.
@@
We now use Theorem A.1 to obtain the sample complexity.
Lemma A.6.
Let $\epsilon\in\left(0,1\right)$, and let $\tau\in\mathbb{N}$ and $\mathcal{H}_{\tau}$ such that $h^{\star}\in\mathcal{H}_{\tau}$. Then for any interpolating model $h_{\mathcal{S}}\in\mathcal{H}_{\tau}$, and $N\geq\frac{\log\left({\tau}/{\delta_{\mathcal{S}}}\right)}{\epsilon}$
$\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N}}\left(\mathcal{L}_{\mathcal{D}}\left(h_{\mathcal{S}}\right)\leq\epsilon\right)\geq 1-\delta_{\mathcal{S}}\,,$
@@
If $h^{\star}\in\mathcal{H}_{\tau}$ then $\mathcal{H}_{\tau}$ is realizable, so we get from Theorem A.1 with
$N\geq\frac{\log\left(\frac{{\left\lvert{\mathcal{H}_{\tau}}\right\rvert}}{\delta_{\mathcal{S}}}\right)}{\epsilon}=\frac{\log\left(\frac{\tau}{\delta_{\mathcal{S}}}\right)}{\epsilon}$
@@
and therefore combining (19), (18) and (16) together with 13 we finish the proof:
@@
We are now ready to prove Theorem A.2:
@@
(of Theorem A.2) Note that $h\sim\mathcal{P}_{\mathcal{S}}$ is interpolating by definition, and therefore, from Lemma A.6, we have that for any $\tau\in\mathbb{N}$ it holds that
$\forall\,\epsilon\geq\frac{\log\left(\frac{\tau}{\delta_{\mathcal{S}}}\right)}{N}\quad\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N},h\sim\mathcal{P}_{\mathcal{S}}}\left(\mathcal{L}_{\mathcal{D}}\left(h\right)\leq\epsilon\middle|h^{\star}\in\mathcal{H}_{\tau}\right)=\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N}}\left(\mathcal{L}_{\mathcal{D}}\left(\bar{h}(S)\right)\leq\epsilon\right)\geq 1-\delta_{\mathcal{S}}\,,$ (20)
Where $\bar{h}\left(S\right)\in\mathcal{H}_{\tau}$ is some arbitrary interpolate, since from Lemma A.6 this probability is equal for any interpolator. For $\tau=\left\lceil\frac{\log\left(\delta_{h}\right)}{\log\left(1-\tilde{p}\right)}\right\rceil$ we similarly get
$\mathbb{P}_{S\sim\mathcal{D}^{N},h\sim\mathcal{P}_{\mathcal{S}}}\left(\mathcal{L}_{\mathcal{D}}(h)\leq\epsilon\right)$ $=$
@@
$\,+\mathbb{P}_{S\sim\mathcal{D}^{N},h\sim\mathcal{P}_{\mathcal{S}}}\left(\mathcal{L}_{\mathcal{D}}(h)\leq\epsilon\middle|h^{\star}\notin\mathcal{H}_{\tau}\right)\mathbb{P}_{h\sim\mathcal{P}_{\mathcal{S}}}\left(h^{\star}\notin\mathcal{H}_{\tau}\right)$
@@
And as before, this is true for any $N\geq\frac{\log\left(\frac{\tau}{\delta_{\mathcal{S}}}\right)}{\epsilon}$. specifically, by Lemma A.7, it is true for $N$ greater than
$\frac{\log\left(\frac{1}{\delta_{\mathcal{S}}}\right)+\log\left(\frac{1}{\tilde{p}}\right)+\log\left(\log\left(\frac{1}{\delta_{h}}\right)\right)+\frac{\tilde{p}}{\log\left(\frac{1}{\delta_{h}}\right)}}{\epsilon}\,,$
@@
$\frac{\log\left(\frac{2}{\delta}\right)+\log\left(\frac{1}{\tilde{p}}\right)+\log\left(\log\left(\frac{2}{\delta}\right)\right)+\frac{\tilde{p}}{\log\left(\frac{2}{\delta}\right)}}{\epsilon}\,,$
@@
$\mathbb{P}_{S\sim\mathcal{D}^{N},h\sim\mathcal{P}_{\mathcal{S}}}\left(\mathcal{L}_{\mathcal{D}}(h)\leq\epsilon\right)$ $\geq\left(1-\frac{\delta}{2}\right)^{2}$
@@
$\left(-\log\left({\tilde{p}}\right)+3\log\left(\frac{2}{\delta}\right)\right)\frac{1}{\epsilon}\,.$
A.3 Proof for the Volume of Generalizing Interpolators (Corollary 3.3)
@@
Corollary A.8 (volume of generalizing interpolators restated).
For $\epsilon,\delta$ as above, and any $N$ larger than
$\frac{-\log\left({\tilde{p}}\right)+6\log\left(\frac{2}{\delta}\right)}{\epsilon}\,,$
@@
$\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N}}\left(\mathbb{P}_{h\sim\mathcal{P_{S}}}\left(\mathcal{L_{D}}\left(h\right)\geq\epsilon\right)\geq\delta\right)\leq\delta\,.$
@@
Using Theorem 3.2, for any
$N\geq\frac{-\log\left({\tilde{p}}\right)+6\log\left(\frac{2}{\delta}\right)}{\epsilon}\geq\frac{-\log\left({\tilde{p}}\right)+3\log\left(\frac{2}{\delta^{2}}\right)}{\epsilon}$
@@
$\delta^{2}$ $\geq\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N},h\sim\mathcal{P_{S}}}\left(\mathcal{L_{D}}\left(h\right)\geq\epsilon\right)$
$=\mathbb{E}_{\mathcal{S}\sim\mathcal{D}^{N},h\sim\mathcal{P_{S}}}\left[\mathbb{I}\left\{\mathcal{L_{D}}\left(h\right)\geq\epsilon\right\}\right]$
$=\mathbb{E}_{\mathcal{S}\sim\mathcal{D}^{N}}\left[\mathbb{E}_{h\sim\mathcal{P_{S}}}\left[\mathbb{\mathbb{I}}\left\{\mathcal{L_{D}}\left(h\right)\geq\epsilon\right\}\right]\right]$
$=\mathbb{E}_{\mathcal{S}\sim\mathcal{D}^{N}}\left[\mathbb{P}_{h\sim\mathcal{P_{S}}}\left(\mathcal{L_{D}}\left(h\right)\geq\epsilon\right)\right]$
@@
$\mathbb{P}_{\mathcal{S}}\left(\mathbb{P}_{h\sim\mathcal{P_{S}}}\left(\mathcal{L_{D}}\left(h\right)\geq\epsilon\right)\geq\delta\right)\leq\frac{\mathbb{E}_{\mathcal{S}\sim\mathcal{D}^{N}}\left[\mathbb{P}_{h\sim\mathcal{P_{S}}}\left(\mathcal{L_{D}}\left(h\right)\geq\epsilon\right)\right]}{\delta}\leq\delta\,,$
@@
$\mathbb{P}_{\mathcal{S}}\left(\mathbb{P}_{h\sim\mathcal{P_{S}}}\left(\mathcal{L_{D}}\left(h\right)<\epsilon\right)\geq 1-\delta\right)\geq 1-\delta\,.$
A.4 Relationship to PAC Bayes
As stated in Section 3, a similar result to Theorem 3.2 could be derived with PAC-Bayes analysis [McAllester, 1999], which typically focuses on the expected population error $\mathbb{E}_{h\sim\mathcal{P}_{\mathcal{S}}}\left[\mathcal{L}_{\mathcal{D}}(h)\right]$ of a sample from the posterior. A standard PAC-Bayes bound [Langford and Seeger, 2002, McAllester, 2003] yields the following result:
Proposition A.9.
Let $\epsilon>0$ and $\delta\in\left(0,1\right)$. Then with sample complexity

### v2

@@ Appendix B Generalization Results
  • A restatement and discussion on the connection to a well-known result from (Shalev-Shwartz & Ben-David, 2014, Corollary 2.3) in Appendix B.1.
  • The proof of our primary generalization result (Lemma 3.2) in Appendix B.2.
  • The proof of Corollary 3.3 in Appendix B.3.
  • An extension of Lemma 3.2 to non-interpolating solutions in Appendix B.4.
  • A further discussion on the relation of Lemma 3.2 to PAC-Bayes is brought in Appendix B.5
  • In Appendix B.6 we prove a refined version of Lemma 3.2 and discuss its implications.
  • The proof of an alternative generalization result, used for proving generalization in the continuous setting, in Appendix B.7.
B.1 Finite Hypothesis Class PAC Generalization Bound
Here we present a classic generalization bound for finite hypothesis class adapted from (Shalev-Shwartz & Ben-David, 2014), which we will use throughout our paper.
Theorem B.1.
[[[[adapted from (Shalev-Shwartz & Ben-David, 2014, Corollary 2.3)]]]] Let $\mathcal{H}$ be a finite hypothesis class. Let $\delta\in\left(0,1\right)$ and $\varepsilon>0$ and let $N$ be an integer that satisfies
$N\geq\frac{\log\left({\left\lvert{\mathcal{H}}\right\rvert}/\delta\right)}{\varepsilon}\,.$
Then, for any realizable data distribution $\mathcal{D}$, with probability of at least $1-\delta$ over the choice of an i.i.d sample $S$ of size $N$, we have that for every interpolating121212(Shalev-Shwartz & Ben-David, 2014, Corollary 2.3) is more general, but we focus on interpolating hypotheses. hypothesis, $h_{\boldsymbol{\theta}}$, it holds that
$\mathcal{L}_{\mathcal{D}}\left(h_{\mathcal{S}}\right)\leq\varepsilon\,.$
Applying Theorem B.1 to quantized neural networks. It is possible to directly obtain a generalization bound for any quantized model, such as those presented in Section 4. Note that the hypothesis class of quantized models with a finite number of parameters, for example, neural networks with finite width and depth, with $M$ parameters and $Q$ quantization levels is finite and of size $\left|\mathcal{H}\right|=Q^{M}$. Therefore, the sample complexity from Theorem B.1 becomes
$N\geq\frac{M\log\left(Q\right)+\log\left(\frac{1}{\delta}\right)}{\varepsilon}$ (13)
Note that Theorem B.1 is used to prove Lemma 3.2.The sample complexity in (13) is dependent on the student parameters, and therefore much worse than the sample complexities we derive throughout our paper.
B.2 Proving the Generalization of Algorithm 1
In this section, we rely on the existence of the teacher to provide guarantee on the generalization of Algorithm 1. We will first restate Lemma 3.2.
Lemma B.2 (Lemma 3.2 restated).
Let $\varepsilon\in\left(0,1\right)$ and $\delta\in\left(0,\frac{1}{5}\right)$, and assume that $\tilde{p}<\frac{1}{2}$. For any $N$ larger than
$\left(-\log\left({\tilde{p}}\right)+3\log\left(\frac{2}{\delta}\right)\right)\frac{1}{\varepsilon}\,,$
@@ B.2 Proving the Generalization of Algorithm 1
$\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N},h\sim\mathcal{P}_{\mathcal{S}}}\left(\mathcal{L}_{\mathcal{D}}(h)\leq\varepsilon\right)\geq 1-\delta\,.$
Proof Outline. The idea is to show that the probability to sample a TE model within a finite number of steps $\tau$ is large. Conditioning on that event, we treat those models as a realizable finite hypothesis class, obtaining the sample complexity using Theorem B.1, which is then bounded, and after some technical details the theorem is obtained.
@@
We now use Theorem B.1 to obtain the sample complexity.
Lemma B.6.
Let $\varepsilon\in\left(0,1\right)$, and let $\tau\in\mathbb{N}$ and $\mathcal{H}_{\tau}$ such that $h^{\star}\in\mathcal{H}_{\tau}$. Then for any interpolating model $h_{\mathcal{S}}\in\mathcal{H}_{\tau}$, and $N\geq\frac{\log\left({\tau}/{\delta_{\mathcal{S}}}\right)}{\varepsilon}$
$\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N}}\left(\mathcal{L}_{\mathcal{D}}\left(h_{\mathcal{S}}\right)\leq\varepsilon\right)\geq 1-\delta_{\mathcal{S}}\,,$
@@
If $h^{\star}\in\mathcal{H}_{\tau}$ then $\mathcal{H}_{\tau}$ is realizable, so we get from Theorem B.1 with
$N\geq\frac{\log\left(\frac{{\left\lvert{\mathcal{H}_{\tau}}\right\rvert}}{\delta_{\mathcal{S}}}\right)}{\varepsilon}=\frac{\log\left(\frac{\tau}{\delta_{\mathcal{S}}}\right)}{\varepsilon}$
@@
and therefore combining (20), (19) and (17) together with 14 we finish the proof:
@@
We are now ready to prove Theorem B.2:
@@
(of Theorem B.2) Note that $h\sim\mathcal{P}_{\mathcal{S}}$ is interpolating by definition, and therefore, from Lemma B.6, we have that for any $\tau\in\mathbb{N}$ it holds that
$\forall\,\varepsilon\geq\frac{\log\left(\frac{\tau}{\delta_{\mathcal{S}}}\right)}{N}\quad\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N},h\sim\mathcal{P}_{\mathcal{S}}}\left(\mathcal{L}_{\mathcal{D}}\left(h\right)\leq\varepsilon\middle|h^{\star}\in\mathcal{H}_{\tau}\right)=\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N}}\left(\mathcal{L}_{\mathcal{D}}\left(\bar{h}(S)\right)\leq\varepsilon\right)\geq 1-\delta_{\mathcal{S}}\,,$ (21)
Where $\bar{h}\left(S\right)\in\mathcal{H}_{\tau}$ is some arbitrary interpolate, since from Lemma B.6 this probability is equal for any interpolator. For $\tau=\left\lceil\frac{\log\left(\delta_{h}\right)}{\log\left(1-\tilde{p}\right)}\right\rceil$ we similarly get
$\mathbb{P}_{S\sim\mathcal{D}^{N},h\sim\mathcal{P}_{\mathcal{S}}}\left(\mathcal{L}_{\mathcal{D}}(h)\leq\varepsilon\right)$ $=$
@@
$\,+\mathbb{P}_{S\sim\mathcal{D}^{N},h\sim\mathcal{P}_{\mathcal{S}}}\left(\mathcal{L}_{\mathcal{D}}(h)\leq\varepsilon\middle|h^{\star}\notin\mathcal{H}_{\tau}\right)\mathbb{P}_{h\sim\mathcal{P}_{\mathcal{S}}}\left(h^{\star}\notin\mathcal{H}_{\tau}\right)$
@@
And as before, this is true for any $N\geq\frac{\log\left(\frac{\tau}{\delta_{\mathcal{S}}}\right)}{\varepsilon}$. specifically, by Lemma B.7, it is true for $N$ greater than
$\frac{\log\left(\frac{1}{\delta_{\mathcal{S}}}\right)+\log\left(\frac{1}{\tilde{p}}\right)+\log\left(\log\left(\frac{1}{\delta_{h}}\right)\right)+\frac{\tilde{p}}{\log\left(\frac{1}{\delta_{h}}\right)}}{\varepsilon}\,,$
@@
$\frac{\log\left(\frac{2}{\delta}\right)+\log\left(\frac{1}{\tilde{p}}\right)+\log\left(\log\left(\frac{2}{\delta}\right)\right)+\frac{\tilde{p}}{\log\left(\frac{2}{\delta}\right)}}{\varepsilon}\,,$
@@
$\mathbb{P}_{S\sim\mathcal{D}^{N},h\sim\mathcal{P}_{\mathcal{S}}}\left(\mathcal{L}_{\mathcal{D}}(h)\leq\varepsilon\right)$ $\geq\left(1-\frac{\delta}{2}\right)^{2}$
@@
$\left(-\log\left({\tilde{p}}\right)+3\log\left(\frac{2}{\delta}\right)\right)\frac{1}{\varepsilon}\,.$
B.3 Proof for the Volume of Generalizing Interpolators (Corollary 3.3)
@@
Corollary B.8 (volume of generalizing interpolators restated).
For $\varepsilon,\delta$ as above, and any $N$ larger than
$\frac{-\log\left({\tilde{p}}\right)+6\log\left(\frac{2}{\delta}\right)}{\varepsilon}\,,$
@@
$\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N}}\left(\mathbb{P}_{h\sim\mathcal{P_{S}}}\left(\mathcal{L_{D}}\left(h\right)\geq\varepsilon\right)\geq\delta\right)\leq\delta\,.$
@@
Using Lemma 3.2, for any
$N\geq\frac{-\log\left({\tilde{p}}\right)+6\log\left(\frac{2}{\delta}\right)}{\varepsilon}\geq\frac{-\log\left({\tilde{p}}\right)+3\log\left(\frac{2}{\delta^{2}}\right)}{\varepsilon}$
@@
$\delta^{2}$ $\geq\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N},h\sim\mathcal{P_{S}}}\left(\mathcal{L_{D}}\left(h\right)\geq\varepsilon\right)$
$=\mathbb{E}_{\mathcal{S}\sim\mathcal{D}^{N},h\sim\mathcal{P_{S}}}\left[\mathbb{I}\left\{\mathcal{L_{D}}\left(h\right)\geq\varepsilon\right\}\right]$
$=\mathbb{E}_{\mathcal{S}\sim\mathcal{D}^{N}}\left[\mathbb{E}_{h\sim\mathcal{P_{S}}}\left[\mathbb{\mathbb{I}}\left\{\mathcal{L_{D}}\left(h\right)\geq\varepsilon\right\}\right]\right]$
$=\mathbb{E}_{\mathcal{S}\sim\mathcal{D}^{N}}\left[\mathbb{P}_{h\sim\mathcal{P_{S}}}\left(\mathcal{L_{D}}\left(h\right)\geq\varepsilon\right)\right]$
@@
$\mathbb{P}_{\mathcal{S}}\left(\mathbb{P}_{h\sim\mathcal{P_{S}}}\left(\mathcal{L_{D}}\left(h\right)\geq\varepsilon\right)\geq\delta\right)\leq\frac{\mathbb{E}_{\mathcal{S}\sim\mathcal{D}^{N}}\left[\mathbb{P}_{h\sim\mathcal{P_{S}}}\left(\mathcal{L_{D}}\left(h\right)\geq\varepsilon\right)\right]}{\delta}\leq\delta\,,$
@@
$\mathbb{P}_{\mathcal{S}}\left(\mathbb{P}_{h\sim\mathcal{P_{S}}}\left(\mathcal{L_{D}}\left(h\right)<\varepsilon\right)\geq 1-\delta\right)\geq 1-\delta\,.$
B.4 Extension to Non-Interpolators
Let $\varepsilon>0$. Assume that there exists a narrow teacher NN $h^{\star}$ s.t $\mathcal{L_{D}}\left(h^{\star}\right)=\varepsilon^{\star}>0$. Denote by $\tilde{p}$ the teacher equivalence probability, and define the posterior distribution
$\mathcal{P_{S}}(h)=\mathcal{P}_{\mathcal{S}}^{\varepsilon}(h)=\mathbb{P}\left(h\,\middle|\,\mathcal{L_{S}}(h)\leq\varepsilon^{\star}+\varepsilon\right)\,.$
In the G&C formulation, sampling from $\mathcal{P}_{\mathcal{S}}^{\varepsilon}$ is equivalent to stopping at the first model satisfying $\mathcal{L_{S}}\left(h\right)\leq\varepsilon^{\star}+\varepsilon$.
Theorem B.9 (Hoeffding’s Inequality).
Let $X_{1},\dots,X_{N}$ be i.i.d random variables with $\mathbb{E}X_{i}=\mu$ and $0\leq X_{i}\leq 1$ a.s. Then for all $t>0$
$\mathbb{P}\left(\frac{1}{N}\sum_{i=1}^{N}X_{i}-\mu\geq t\right)\leq\exp\left(-2Nt^{2}\right)$
and
$\mathbb{P}\left(\left|\frac{1}{N}\sum_{i=1}^{N}X_{i}-\mu\right|\geq t\right)\leq 2\exp\left(-2Nt^{2}\right)\,.$
By definition, $\mathcal{L_{S}}\left(h\right)=\frac{1}{N}\sum_{\left(\mathbf{x},y\right)\in\mathcal{S}}\mathbb{I}\left\{h\left(\mathbf{x}\right)\neq y\right\}$ so using Hoeffding’s inequality
$\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N}}\left(\mathcal{L_{S}}\left(h^{\star}\right)\geq\varepsilon^{\star}+\varepsilon\right)\leq e^{-2\varepsilon^{2}N}\,.$
Lemma B.10.
Let $\varepsilon\in\left(0,\frac{1}{2}-\varepsilon^{\star}\right)$, and let $\tau\in\mathbb{N}$, and $\mathcal{H}_{\tau}$ s.t $h^{\star}\in\mathcal{H}_{\tau}$. Then for any model $\tilde{h}\in\mathcal{H}_{\tau}$ satisfying $\mathcal{L_{S}}\left(\tilde{h}\right)\leq\varepsilon^{\star}+\varepsilon$
$\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N}}\left(\mathcal{L_{D}}\left(\tilde{h}\right)\leq\varepsilon^{\star}+2\varepsilon,\mathcal{L_{S}}\left(h^{\star}\right)\leq\varepsilon^{\star}+\varepsilon\right)\geq 1-\left(2\tau+1\right)e^{-2\varepsilon^{2}N}\,.$
Proof.
Part 1: We will show that
$\mathcal{L_{D}}\left(\tilde{h}\right)-\varepsilon^{\star}\leq\max_{h\in\mathcal{H}_{\tau}}\left|\mathcal{L_{D}}\left(h\right)-\mathcal{L_{S}}\left(h\right)\right|+\varepsilon\,.$
Since $h^{\star}\in\mathcal{H}_{\tau}$, w.p. $1-e^{-2\varepsilon^{2}N}$ it holds that $\mathcal{L_{S}}\left(h^{\star}\right)\leq\varepsilon^{\star}+\varepsilon$ so there exists such $\tilde{h}\in\mathcal{H}_{\tau}$. Then
$\mathcal{L_{D}}\left(\tilde{h}\right)-\varepsilon^{\star}$ $=\mathcal{L_{D}}\left(\tilde{h}\right)-\mathcal{L_{S}}\left(\tilde{h}\right)+\mathcal{L_{S}}\left(\tilde{h}\right)-\varepsilon^{\star}$
$\leq\left(\mathcal{L_{D}}\left(\tilde{h}\right)-\mathcal{L_{S}}\left(\tilde{h}\right)\right)+\varepsilon^{\star}+\varepsilon-\varepsilon^{\star}$
$\leq\max_{h\in\mathcal{H}_{\tau}}\left|\mathcal{L_{D}}\left(h\right)-\mathcal{L_{S}}\left(h\right)\right|+\varepsilon\,.$
Part 2: Using the union bound and then Hoeffding’s inequality
$\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N}}\left(\mathcal{L_{D}}\left(\tilde{h}\right)-\varepsilon^{\star}>2\varepsilon\right)$ $\leq\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N}}\left(\max_{h\in\mathcal{H}_{\tau}}\left|\mathcal{L_{D}}\left(h\right)-\mathcal{L_{S}}\left(h\right)\right|+\varepsilon>2\varepsilon\right)$
$=\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N}}\left(\max_{h\in\mathcal{H}_{\tau}}\left|\mathcal{L_{D}}\left(h\right)-\mathcal{L_{S}}\left(h\right)\right|>\varepsilon\right)$
$=\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N}}\left(\exists h\in\mathcal{H}_{\tau}\,:\,\left|\mathcal{L_{D}}\left(h\right)-\mathcal{L_{S}}\left(h\right)\right|>\varepsilon\right)$
$\leq\sum_{h\in\mathcal{H}_{\tau}}\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N}}\left(\left|\mathcal{L_{D}}\left(h\right)-\mathcal{L_{S}}\left(h\right)\right|>\varepsilon\right)$
$\leq\sum_{h\in\mathcal{H}_{\tau}}2\exp\left(-2N\varepsilon^{2}\right)$
$=2\tau e^{-2N\varepsilon^{2}}\,.$
Part 3: Combining the probability lower bounds using the union bound
$\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N}}\left(\mathcal{L_{D}}\left(\tilde{h}\right)\leq\varepsilon^{\star}+2\varepsilon,\mathcal{L_{S}}\left(h^{\star}\right)\leq\varepsilon^{\star}+\varepsilon\right)$ $\geq 1-2\tau e^{-2\varepsilon^{2}N}-e^{-2\varepsilon^{2}N}$
$=1-\left(2\tau+1\right)e^{-2\varepsilon^{2}N}\,.$
With Lemma B.5 and Lemma B.10 we deduce the following.
Theorem B.11 (Generalization of non-interpolating student).